OpenStudy (anonymous):

proove that if x^2 is divisible by 6 then x is divisible by 6

OpenStudy (anonymous):

we can write x=6y, where y is an integer (as x is divisible by 6) so x^2=36y^2 can you do the rest?

OpenStudy (anonymous):

it's the other way around, prove that if x^2=6y, then x can be written as x= 6w. x,y,w are all natural numbers

OpenStudy (zarkon):

prove the contapositive

OpenStudy (anonymous):

I knew the idea of proving it is by contapositive, but im having trouble doing it

OpenStudy (anonymous):

oh okay, my bad. so all the prime factors should appear in multiple of 2. for example, if x=x1^a * x2^b * x3^c... where x1, x2, x3.. are prime factors of x then x^2= x1^2a * x2 ^2b... notice that all the prime factors appear 2a, 2b, 2c... times in x^2, that means they appear in multiple of 2 now x^2=6y x^2=2*3*y 2 and 3 should appear at least two times each as the number is a square so, x^2=2*3*2*3*y1 where y1=y/6 now y1 also has to be a square number so x=6*sqrt (y1) or, x=6 * y2 (y2=sqrt y1, y1 is a square number, so y2 is a whole number) so x is divisible by 6 when x^2 is divisible by 6

OpenStudy (anonymous):

so if i wanna prove it by contapositive i say : if x is not divisible by 6 then x^2 is not divisible by 6 \[\forall k \in \mathbb{N}, x \neq 6 k \] \[\therefore \forall w \in \mathbb{N} , x ^{2} \neq 6 w \]

OpenStudy (anonymous):

is that enough of a proof ?

OpenStudy (anonymous):

contrapositive means assume x is divisible by 6 but x^2 is not divisible by 6, prove the opposite

OpenStudy (anonymous):

retaining the assumption x is divisible by 6

OpenStudy (anonymous):

are you sure? coz my text book indicate the contrapositive of \[A =>B\] is \[-B => -A\]

OpenStudy (anonymous):

if you prove if x is not divisible by 6 implies x^2 is not divisible by 6, does it necessarily prove if x^2 is divisible by 6 then x is divisible by 6 ? i am not sure. but at least i am sure about the way i solved

ganeshie8 (ganeshie8):

the direct proof works, and the contrapositive also works. contrapositive follows the truth value of original statement

OpenStudy (anonymous):

Yeah.

ganeshie8 (ganeshie8):

\(\large A \implies B\) and \(\large \neg B \implies \neg A\) are equivalent statements

OpenStudy (anonymous):

So my question now is, how do you get from \[x \neq 6 k \] to \[x^2 \neq 6 w\]

ganeshie8 (ganeshie8):

\[\forall k \in \mathbb{N}, x \neq 6 k \] \[\therefore \forall k \in \mathbb{N} , x ^{2} \neq (6 k)^2 \] \[\therefore \forall w \in \mathbb{N} , x ^{2} \neq 6 w \]

ganeshie8 (ganeshie8):

i don't see anything to add here, that should end the proof ^

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yw!